# Dist Loss: Enhancing Regression in Few-Shot Region through Distribution Distance Constraint

### v1

@@ 4.1 Benchmarks and baselines
We evaluated our method on three datasets, focusing on tasks of age estimation and potassium concentration prediction. The IMDI-WIKI-DIR dataset Yang et al. (2021), derived from the IMDB-WIKI dataset Rothe et al. (2018), consists of 213,553 facial image pairs annotated with age information. This dataset is partitioned into 191,509 samples for training, 11,022 for validation, and 11,022 for testing. The AgeDB-DIR dataset Yang et al. (2021), derived from the AgeDB dataset Moschoglou et al. (2017), comprises 16,488 facial image pairs with age annotations. It is divided into 12,208 samples for training, 2,140 for validation, and 2,140 for testing. The ECG-Ka-DIR dataset, sourced from the MIMIC-IV dataset Johnson et al. (2020), includes 375,745 pairs of single-lead ECG signals paired with potassium concentration values. This dataset is divided into 365,549 samples for training, 5,098 for validation, and 5,098 for testing. All these datasets are characterized by imbalanced training sets and balanced validation and test sets. The label distributions of these three datasets are shown in Figure 3. Please refer to Appendix A.1 and A.2 for baseline and implementation details.
@@
(c) ECG-Ka-DIR
Figure 3: Overview of label distributions in the training sets for the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-Ka-DIR datasets. The classification of shot types for IMDB-WIKI-DIR and AgeDB-DIR follows the definitions provided in Yang et al. (2021).
@@ 4.2 Evaluation metrics
Following the evaluation metrics of Yang et al. (2021), we report the results for four shots: all, many, median, and few, where all represents the entire dataset, and many/median/few correspond to areas of high/medium/low sample density within the dataset. For the IMDB-WIKI-IR and AgeDB-DIR datasets, we maintain consistency with previous studies, where few/median/many correspond to areas with fewer than 20, between 20-100, and more than 100 samples, respectively. For the ECG-Ka-DIR dataset, assuming that the maximum number of samples for a single label is $n_{\text{max}}$, we define areas with more than 0.5 $n_{\text{max}}$, between 0.15-0.5 $n_{\text{max}}$, and fewer than 0.15 $n_{\text{max}}$ samples as many/median/few shots areas, respectively. For each dataset, we report the mean absolute error (MAE) and the geometric mean (GM).
@@ 4.3 Main results
Table 1 presents the results of baselines and our method in the few-shot region across three datasets, along with a comparison of these results. For detailed results on each dataset, please refer to Appendix A.3. This table is divided into two sections. The first section displays the results of the baselines and our method, with the best results highlighted in bold and red. The second section shows the improvement of our method over each baseline, with green bold indicating superior performance of our method and blue bold indicating otherwise. From the first section of the table, it is evident that our method achieves the best results in five out of six metrics across the three datasets, with SOTA performances of 22.550, 9.122, and 1.329 on the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-Ka-DIR datasets, respectively. The second section reveals that our method outperforms in 28 out of 30 metrics. Notably, compared to Balanced MSE, which also involves fine-tuning the linear layers of a pre-trained model and employs data distribution priors, our method demonstrates superior performance in the few-shot region, highlighting the effectiveness of our approach.
Table 2 further illustrates the complementary nature of our method with existing approaches. This table is divided into five sections, each showcasing the results of one baseline and the combined results with our method, with the best results within each section highlighted in bold and black. From this table, it is shown that our method achieves better results in 26 out of 30 metrics. Taking the MAE metric as an example, incorporating our method leads to improved performance in the few-shot region across all three datasets, achieving the best results of 22.331, 9.110, and 1.325 on the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-Ka-DIR datasets, respectively. These experimental results demonstrate a key advantage of our method, namely its ability to effectively complement existing methods, thereby enhancing model performance in the few-shot region.
Table 1: Results are presented for the few-shot region on the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-Ka-DIR datasets. The first section of the table reports the results of baselines and our method, with the best results highlighted in bold and red. In the second section, improvements over corresponding baselines are reported in bold and green, while decreases in performance are reported in bold and blue.
@@
IMDB-WIKI-DIR AgeDB-DIR ECG-Ka-DIR IMDB-WIKI-DIR AgeDB-DIR ECG-Ka-DIR
@@
Table 2: Results are presented for the few-shot region on the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-Ka-DIR datasets. Each section of the table reports the results of a baseline and the baseline incorporating our method, with the better results highlighted in bold.
@@
IMDB-WIKI-DIR AgeDB-DIR ECG-Ka-DIR IMDB-WIKI-DIR AgeDB-DIR ECG-Ka-DIR
@@
Table 3: Time consumption (in seconds) of one training epoch for the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-Ka-DIR datasets, with batch sizes of 64, 64, and 256, respectively.
IMDB-WIKI-DIR AgeDB-DIR ECG-Ka-DIR
@@ 4.4 Time consumption Analysis
Table 3 presents the time required to train each method for one epoch on the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-Ka-DIR datasets, with all times reported in seconds. It can be observed that Balanced MSE and Dist Loss have the shortest training times, attributed to their approach of fine-tuning the model’s linear layers. The time consumption of LDS and the vanilla model are largely consistent, as these methods only weight the loss function without significantly increasing computational load. For methods operating at the feature level, including FDS, Ranksim, and ConR, a notable increase in model training time is evident, due to the computational intensity associated with feature-level operations.
@@ 4.5.1 Different loss functions for sequence difference measurement
Dist Loss employs the loss function $L(\cdot)$ to measure the difference between two sequences. In this ablation study, we demonstrate the effects of using different functions, considering the probability-based inversely weighted MAE and MSE losses. The experimental results are shown in Table 4, where detailed results on each dataset are shown in Appendix A.4.2. The table illustrates that Dist Loss reliably boosts model accuracy in the few-shot region.
Table 4: Ablation study on loss functions measuring sequence difference. $L_{1}$ represents MAE Loss, $L_{2}$ represents MSE Loss, $INV-$ denotes the probability-based inversely weighted version of these loss functions. Results on the few-shot region are reported, with the best results in each section are in bold.
@@
IMDB-WIKI-DIR AgeDB-DIR ECG-Ka-DIR IMDB-WIKI-DIR AgeDB-DIR ECG-Ka-DIR
@@ 4.5.2 Different batch sizes for distribution distance approximation
Dist Loss estimates the overall distribution distance between predictions and labels by measuring batch-wise distances during training. This ablation study evaluates the sensitivity of model accuracy to batch size, as detailed in Table 5. We examined batch sizes of 256, 512, and 768, adopting 256 as a standard based on prior research Yang et al. (2021); Gong et al. (2022). The findings show negligible variations in performance with different batch sizes. This could be attributed to the fact that accurate distribution information is more critical during sampling than the precise accuracy of individual pseudo-labels.
@@ 4.5.3 Dist Loss Surpasses Existing Methods in the Median-Shot Region
As depicted in the supplementary Tables 6, 7, and 8 within Appendix A.3, Dist Loss delivers SOTA results, excelling not only in few-shot regions but also in median-shot regions. In our comparison with current methods, Dist Loss achieved the lowest MAE and the second-lowest GM on the IMDB-WIKI-DIR and AgeDB-DIR datasets, with scores of 12.614/7.686 and 7.315/4.563, respectively. Similarly, on the ECG-Ka-DIR dataset, it secured the highest GM and the second-lowest MAE, recording 0.445 and 0.674, respectively. Moreover, our experiments show that integrating Dist Loss with existing methods consistently improved performance in median-shot regions when measured by both MAE and GM, surpassing the results of using those methods alone on IMDB-WIKI-DIR and AgeDB-DIR datasets. On the ECG-Ka-DIR dataset, this integration notably increased the GM. In conclusion, these findings validate Dist Loss’s efficacy in enhancing model accuracy in both few-shot and median-shot regions.
@@ 5 Conclusion
In this study, we address the significant escalation of prediction errors in few-shot regions, a prevalent challenge in deep imbalanced regression. By leveraging distribution priors, we introduce a novel loss function, Dist Loss, designed to align the model’s prediction distribution with the label distribution throughout the training process. Our extensive experimental evaluation demonstrates that Dist Loss effectively enhances prediction accuracy in few-shot regions, achieving state-of-the-art performance. Furthermore, our results indicate that Dist Loss can be seamlessly integrated with existing methods to further augment their efficacy. We hope our work underscores the critical role of integrating distribution information in tackling deep imbalanced regression tasks.
@@ A.1 Baselines
To ensure a fair comparison, we followed the experimental setup of Yang et al. (2021) on the IMDB-WIKI-DIR and AgeDB-DIR datasets, i.e., using ResNet-50 as the network architecture and training for 90 epochs. For the ECG-Ka-DIR dataset, we employed the ResNet variant Net1D Hong et al. (2020) as the network architecture. Given that previous work Yang et al. (2021); Ren et al. (2022); Gong et al. (2022); Keramati et al. (2024) has demonstrated superior performance over loss reweighting and regressor re-training (RRT) in deep imbalanced regression tasks, we do not include these methods as baselines in this paper. Instead, we focused on widely recognized approaches in the field: LDS, FDS Yang et al. (2021), Ranksim Gong et al. (2022), ConR Keramati et al. (2024), and Balanced MSE Ren et al. (2022). LDS and FDS encourage local similarities in label and feature space, while Ranksim and ConR leverage contrastive learning to translate label similarities into the feature space. Balanced MSE, based on label distribution priors, restores a balanced distribution from an imbalanced dataset. Our experimental findings indicate that not only does our method achieve SOTA performance in few-shot regions, but it also enhances existing methods, offering a complementary strategy to boost their efficacy.
@@ A.2 Implementation Details
We trained all models on the IMDB-WIKI-DIR and AgeDB-DIR datasets using a single NVIDIA GeForce RTX 3090 GPU and on the ECG-Ka-DIR dataset using a single NVIDIA GeForce RTX 4090 GPU. To ensure a fair comparison, we followed the training, validation, and test set divisions from Yang et al. (2021) for the IMDB-WIKI-DIR and AgeDB datasets. During training with Dist Loss, we used the same strategy as Balanced MSE, fine-tuning the linear layer based on pre-trained model (vanilla model) parameters. This approach integrates our method with existing methods, using their model parameters as the starting point for fine-tuning. Additionally, we used probability-based inversely weighted MSE to measure sequence difference in Dist Loss for all datasets, setting the distribution loss component weight to 1.
@@ A.2.1 IMDB-WIKI-DIR
On the IMDB-WIKI-DIR dataset, we selected ResNet-50 as the network architecture. During training, the training epochs were set to 90, with an initial learning rate of 0.001, which was reduced to 1/10 of its value at the 60th and 80th epochs. We employed the Adam optimizer with a momentum of 0.9 and a weight decay of 0.0001. For our method and Balanced MSE, we used a batch size of 512. For the other baselines, we followed the experimental setups from their original papers. It should be noted that the original training epochs for Ranksim and ConR in their respective papers were 120, which we adjusted to 90 in our experiments to ensure a fair comparison.
@@ A.2.2 AgeDB-DIR
A.2.3 ECG-Ka-DIR
On the ECG-Ka-DIR dataset, we utilized the ResNet variant, Net1D Hong et al. (2020), as our network architecture. The training was set for 10 epochs with an initial learning rate of 0.001, which was reduced to 1/10 of its initial value at the 5th and 8th epochs. We employed the Adam optimizer with a momentum of 0.9 and a weight decay of 0.00001. A batch size of 512 was used for all methods. Additionally, for ConR, we constructed positive and negative sample pairs by adding Gaussian noise.
@@ A.2.3 ECG-Ka-DIR
Table 8: Comprehensive results on the ECG-Ka-DIR dataset are presented. The table highlights the best results in each section using bold font. Additionally, the best result in each column is indicated in bold and red.
@@ A.3 Comprehensive experimental results
Tables 6, 7, and 8 present a comprehensive overview of our experimental results on the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-Ka-DIR datasets. The results indicate that our method achieves improvements in model performance on median-shot and few-shot regions without compromising overall error rates. This further demonstrates the effectiveness of our method in sparse data regions.
@@ A.4.1 Different batch sizes for distribution distance approximation
Table 9 illustrates the impact of varying batch sizes on the final performance across three datasets: IMDB-WIKI-DIR, AgeDB-DIR, and ECG-Ka-DIR. The results indicate that there is no significant difference in performance among different batch sizes. This observation suggests that the generation of pseudo-labels primarily requires an approximation of the distribution information, rather than the precise accuracy of every individual label value.
Table 9: Ablation study examining the impact of batch size on model performance across the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-Ka-DIR datasets.
@@
Tables 10, 11, and 12 present the comprehensive results of using different loss functions on IMDB-WIKI-DIR, AgeDB-DIR, and ECG-Ka-DIR, respectively. It is evident that existing methods, when augmented with Dist Loss, demonstrate superior performance on samples within few-shot regions.
@@
Table 12: An ablation study on loss functions on the ECG-Ka-DIR dataset. $L_{1}$ represents MAE Loss, $L_{2}$ represents MSE Loss, $INV-$ denotes the probability-based inversely weighted version of these loss functions. Results on the few-shot region are reported.

### v2

@@ 4.1 Benchmarks and baselines
We evaluated our method on three datasets, focusing on tasks of age estimation and potassium concentration prediction. The IMDI-WIKI-DIR dataset Yang et al. (2021), derived from the IMDB-WIKI dataset Rothe et al. (2018), consists of 213,553 facial image pairs annotated with age information. This dataset is partitioned into 191,509 samples for training, 11,022 for validation, and 11,022 for testing. The AgeDB-DIR dataset Yang et al. (2021), derived from the AgeDB dataset Moschoglou et al. (2017), comprises 16,488 facial image pairs with age annotations. It is divided into 12,208 samples for training, 2,140 for validation, and 2,140 for testing. The ECG-K-DIR dataset, sourced from the MIMIC-IV dataset Johnson et al. (2020), includes 375,745 pairs of single-lead ECG signals paired with potassium concentration values. This dataset is divided into 365,549 samples for training, 5,098 for validation, and 5,098 for testing. All these datasets are characterized by imbalanced training sets and balanced validation and test sets. The label distributions of these three datasets are shown in Figure 4. Please refer to Appendix A.1 and A.2 for baseline and implementation details.
@@
(c) ECG-K-DIR
Figure 4: Overview of label distributions in the training sets for the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-K-DIR datasets. The classification of shot types for IMDB-WIKI-DIR and AgeDB-DIR follows the definitions provided in Yang et al. (2021).
@@ 4.2 Evaluation metrics
Following the evaluation metrics of Yang et al. (2021), we report the results for four shots: all, many, median, and few, where all represents the entire dataset, and many/median/few correspond to areas of high/medium/low sample density within the dataset. For the IMDB-WIKI-IR and AgeDB-DIR datasets, we maintain consistency with previous studies, where few/median/many correspond to areas with fewer than 20, between 20-100, and more than 100 samples, respectively. For the ECG-K-DIR dataset, assuming that the maximum number of samples for a single label is $n_{\text{max}}$, we define areas with more than 0.5 $n_{\text{max}}$, between 0.15-0.5 $n_{\text{max}}$, and fewer than 0.15 $n_{\text{max}}$ samples as many/median/few shots areas, respectively. For each dataset, we report the mean absolute error (MAE) and the geometric mean (GM).
@@ 4.3 Main results
Table 1 presents the results of baselines and our method in the few-shot region across three datasets, along with a comparison of these results. For detailed results on each dataset, please refer to Appendix A.3. This table is divided into two sections. The first section displays the results of the baselines and our method, with the best results highlighted in bold and red. The second section shows the improvement of our method over each baseline, with green bold indicating superior performance of our method and blue bold indicating otherwise. From the first section of the table, it is evident that our method achieves the best results in five out of six metrics across the three datasets, with SOTA performances of 22.550, 9.122, and 1.329 on the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-K-DIR datasets, respectively. The second section reveals that our method outperforms in 28 out of 30 metrics. Notably, compared to Balanced MSE, which also involves fine-tuning the linear layers of a pre-trained model and employs data distribution priors, our method demonstrates superior performance in the few-shot region, highlighting the effectiveness of our approach.
Table 2 further illustrates the complementary nature of our method with existing approaches. This table is divided into five sections, each showcasing the results of one baseline and the combined results with our method, with the best results within each section highlighted in bold and black. From this table, it is shown that our method achieves better results in 26 out of 30 metrics. Taking the MAE metric as an example, incorporating our method leads to improved performance in the few-shot region across all three datasets, achieving the best results of 22.331, 9.110, and 1.325 on the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-K-DIR datasets, respectively. These experimental results demonstrate a key advantage of our method, namely its ability to effectively complement existing methods, thereby enhancing model performance in the few-shot region.
Table 1: Results are presented for the few-shot region on the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-K-DIR datasets. The first section of the table reports the results of baselines and our method, with the best results highlighted in bold and red. In the second section, improvements over corresponding baselines are reported in bold and green, while decreases in performance are reported in bold and blue.
@@
IMDB-WIKI-DIR AgeDB-DIR ECG-K-DIR IMDB-WIKI-DIR AgeDB-DIR ECG-K-DIR
@@
Table 2: Results are presented for the few-shot region on the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-K-DIR datasets. Each section of the table reports the results of a baseline and the baseline incorporating our method, with the better results highlighted in bold.
@@
IMDB-WIKI-DIR AgeDB-DIR ECG-K-DIR IMDB-WIKI-DIR AgeDB-DIR ECG-K-DIR
@@
Table 3: Time consumption (in seconds) of one training epoch for the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-K-DIR datasets, with batch sizes of 64, 64, and 256, respectively.
IMDB-WIKI-DIR AgeDB-DIR ECG-K-DIR
@@ 4.4 Time consumption Analysis
Table 3 presents the time required to train each method for one epoch on the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-K-DIR datasets, with all times reported in seconds. It can be observed that Balanced MSE and Dist Loss have the shortest training times, attributed to their approach of fine-tuning the model’s linear layers. The time consumption of LDS and the vanilla model are largely consistent, as these methods only weight the loss function without significantly increasing computational load. For methods operating at the feature level, including FDS, Ranksim, and ConR, a notable increase in model training time is evident, due to the computational intensity associated with feature-level operations.
@@ 4.5.1 Different loss functions for sequence difference measurement
Dist Loss employs the loss function $L(\cdot)$ to measure the difference between two sequences. In this ablation study, we demonstrate the effects of using different functions, including the inverse probability weighted MAE and MSE losses. The experimental results are shown in Table 4, with detailed results on each dataset provided in Appendix A.4.2. The table illustrates that Dist Loss reliably improves model accuracy in the few-shot region.
Table 4: Ablation study on loss functions measuring sequence difference. $L_{1}$ represents MAE Loss, $L_{2}$ represents MSE Loss, $INV-$ denotes the the version of these loss functions that are inverse probability weighted. Results on the few-shot region are reported, with the best results in each section are in bold.
@@
IMDB-WIKI-DIR AgeDB-DIR ECG-K-DIR IMDB-WIKI-DIR AgeDB-DIR ECG-K-DIR
@@ 4.5.2 Different batch sizes for distribution distance approximation
Dist Loss estimates the overall distribution distance between predictions and labels by measuring batch-wise distances during training. This ablation study evaluates the sensitivity of model accuracy to batch size, as detailed in Table 5. We examined batch sizes of 256, 512, and 768, adopting 256 as a standard based on prior research Yang et al. (2021); Gong et al. (2022). The findings show negligible variations in performance with different batch sizes. This could be attributed to the fact that accurately reflecting the distribution information during sampling is more important than requiring each value in the generated pseudo-sequence to be perfectly accurate, which in turn allows for a smaller batch size.
@@ 4.5.3 Dist Loss Surpasses Existing Methods in the Median-Shot Region
As depicted in the supplementary Tables 6, 7, and 8 within Appendix A.3, Dist Loss delivers SOTA results, excelling not only in few-shot regions but also in median-shot regions. In our comparison with current methods, Dist Loss achieved the lowest MAE and the second-lowest GM on the IMDB-WIKI-DIR and AgeDB-DIR datasets, with scores of 12.614/7.686 and 7.315/4.563, respectively. Similarly, on the ECG-K-DIR dataset, it secured the highest GM and the second-lowest MAE, recording 0.445 and 0.674, respectively. Moreover, our experiments show that integrating Dist Loss with existing methods consistently improved performance in median-shot regions when measured by both MAE and GM, surpassing the results of using those methods alone on IMDB-WIKI-DIR and AgeDB-DIR datasets. On the ECG-K-DIR dataset, this integration notably increased the GM. In conclusion, these findings validate Dist Loss’s efficacy in enhancing model accuracy in both few-shot and median-shot regions.
@@ 5 Conclusion
In this study, we address the significant escalation of prediction errors in few-shot regions, a prevalent challenge in imbalanced regression. By leveraging distribution priors, we introduce a novel loss function, Dist Loss, designed to align the model’s prediction distribution with the label distribution throughout the training process. Our extensive experimental evaluation demonstrates that Dist Loss effectively enhances prediction accuracy in few-shot regions, achieving state-of-the-art performance. Furthermore, our results indicate that Dist Loss can be seamlessly integrated with existing methods to further augment their efficacy. We hope our work underscores the critical role of integrating distribution information in tackling deep imbalanced regression tasks.
@@ A.1 Baselines
To ensure a fair comparison, we followed the experimental setup of Yang et al. (2021) on the IMDB-WIKI-DIR and AgeDB-DIR datasets, i.e., using ResNet-50 as the network architecture and training for 90 epochs. For the ECG-K-DIR dataset, we employed the ResNet variant Net1D Hong et al. (2020) as the network architecture. Given that previous work Yang et al. (2021); Ren et al. (2022); Gong et al. (2022); Keramati et al. (2024) has demonstrated superior performance over loss reweighting and RRT in deep imbalanced regression tasks, we do not include these methods as baselines in this paper. Instead, we focused on widely recognized approaches in the field: LDS, FDS Yang et al. (2021), Ranksim Gong et al. (2022), ConR Keramati et al. (2024), and Balanced MSE Ren et al. (2022). LDS and FDS encourage local similarities in label and feature space, while Ranksim and ConR leverage contrastive learning to translate label similarities into the feature space. Balanced MSE, based on label distribution priors, restores a balanced distribution from an imbalanced dataset. Our experimental findings indicate that not only does our method achieve SOTA performance in few-shot regions, but it also enhances existing methods, offering a complementary strategy to boost their efficacy.
@@ A.2 Implementation Details
We trained all models on the IMDB-WIKI-DIR and AgeDB-DIR datasets using a single NVIDIA GeForce RTX 3090 GPU and on the ECG-K-DIR dataset using a single NVIDIA GeForce RTX 4090 GPU. To ensure a fair comparison, we followed the training, validation, and test set divisions from Yang et al. (2021) for the IMDB-WIKI-DIR and AgeDB datasets. During training with Dist Loss, we used the same strategy as Balanced MSE, fine-tuning the linear layer based on pre-trained model (vanilla model) parameters. This approach integrates our method with existing methods, using their model parameters as the starting point for fine-tuning. Additionally, we used inverse probability weighted MSE to measure sequence difference in Dist Loss for all datasets, setting the distribution loss component weight to 1.
@@ A.2.1 IMDB-WIKI-DIR
On the IMDB-WIKI-DIR dataset, we selected ResNet-50 as the network architecture. During training, the training epochs were set to 90, with an initial learning rate of 0.001, which was reduced to 1/10 of its value at the 60th and 80th epochs. We employed the Adam optimizer with a momentum of 0.9 and a weight decay of 0.0001. For our method and Balanced MSE, we used a batch size of 512. For the other baselines, we followed the experimental setups from their original papers. It should be noted that the original training epochs for ConR was 120, which we adjusted to 90 in our experiments to ensure a fair comparison.
@@ A.2.2 AgeDB-DIR
A.2.3 ECG-K-DIR
On the ECG-K-DIR dataset, we utilized the ResNet variant, Net1D Hong et al. (2020), as our network architecture. The training was set for 10 epochs with an initial learning rate of 0.001, which was reduced to 1/10 of its initial value at the 5th and 8th epochs. We employed the Adam optimizer with a momentum of 0.9 and a weight decay of 0.00001. A batch size of 512 was used for all methods. Additionally, for ConR, we constructed positive and negative sample pairs by adding Gaussian noise.
@@ A.2.3 ECG-K-DIR
Table 8: Comprehensive results on the ECG-K-DIR dataset are presented. The table highlights the best results in each section using bold font. Additionally, the best result in each column is indicated in bold and red.
@@ A.3 Comprehensive experimental results
Tables 6, 7, and 8 present a comprehensive overview of our experimental results on the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-K-DIR datasets. The results indicate that our method achieves improvements in model performance on median-shot and few-shot regions without compromising overall error rates. This further demonstrates the effectiveness of our method in sparse data regions.
@@ A.4.1 Different batch sizes for distribution distance approximation
Table 9 illustrates the impact of varying batch sizes on the final performance across three datasets: IMDB-WIKI-DIR, AgeDB-DIR, and ECG-K-DIR. The results indicate that there is no significant difference in performance among different batch sizes. This observation suggests that the generation of pseudo-labels primarily requires an approximation of the distribution information, rather than the precise accuracy of every individual label value.
Table 9: Ablation study examining the impact of batch size on model performance across the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-K-DIR datasets.
@@
Tables 10, 11, and 12 present the comprehensive results of using different loss functions on IMDB-WIKI-DIR, AgeDB-DIR, and ECG-K-DIR, respectively. It is evident that existing methods, when augmented with Dist Loss, demonstrate superior performance on samples within few-shot regions.
@@
Table 12: An ablation study on loss functions on the ECG-K-DIR dataset. $L_{1}$ represents MAE Loss, $L_{2}$ represents MSE Loss, $INV-$ denotes the probability-based inversely weighted version of these loss functions. Results on the few-shot region are reported.